OpenStudy (anonymous):

What is an equation for the line that passes through points (–1, –4) and (1, 4)? Write the equation in slope-intercept form.

OpenStudy (aroub):

slope: 4+4/1+1 8/2 4 equation: y=mx+b , m=4 , (1,4) 4=4(1)+b 4=4+b 0=b => y=4x